quasia:

Theresa earned scores of 88 points, 84 points, 88 points, and 91 points on four tests. What is the lowest score she can get on the fifth test and still finish with an average score of 90 points? (3 points) 96 points 97 points 98 points 99 points

Extrinix:

So, if you find the average of all four scores, you get 87.75. This means that you need a number even higher than 91 to raise the estimated average, so let's try 95. \(\dfrac{88+84+88+91+95}{5}\) This would get you 89.2, close but not enough, now let's try 96. \(\dfrac{88+84+88+91+96}{5}\) This time you get 89.4, notice that it went up by .2, so we can assume that each time we raise the grade by 1 point, the average increases by .2, so we can figure this out by continuing the chain, 89.4 + .2 + .2 + .2 = 90 So by continuing this chain we did this: 96 > 97 > 98 > 99 So your lowest grade to have an average of 90 is a \(99\).

quasia:

so 99 is basically the answer rii ?.

Extrinix:

Yeah.

surjithayer:

for 5 tests average=90 total of 5 scores=90 *5=450 total of 4 scores=88+84+88+91 =351 score in 5th test=450-351=99